Please solve this system completely and show all work. Solve for x, y, and z.
x + 2y + z = 7
3x - y + z = -12
4x + 3y -2z = 9

Answer by sachi(548)

x + 2y + z = 7...............1
3x - y + z = -12..............2
4x + 3y -2z = 9...............3
eqn 2 X 2+eqn 1
x + 2y + z = 7...............1
6x - 2y + 2z = -24..............3
---------------------
7x +3z =-17.............5
eqn 2X 3 +eqn 3
9x -3y +3z = -36..............6
4x + 3y -2z = 9...............3
-----------------------
13x +z=-27..............7
eq5 - eqn 7 X 3
7x +3z =-17.............5
39x +3z=-81..............9
- - +
-----------
-32x=64
or x=-2
from eqn 5
7*(-2)+z=-17
or z=-17+14=-3
from eqn 1
-2 + 2y -3 = 7.
or 2y=7+5=12
or y=6